Solve the equation 1/8 + 5/6x = -2/9

Found 4 solutions by helpnalgebra, stanbon, Alan3354, 111456:
Answer by helpnalgebra(91)

(1)/(8)+(5)/(6)*x=-(2)/(9)
(1)/(8)+(5x)/(6)=-(2)/(9)
(5x)/(6)=-(1)/(8)-(2)/(9)
(5x)/(6)=-(25)/(72)
(5x)/(6)*6=-(25)/(72)*6
5x=-(25)/(72)*6
5x=-(25)/(12)
(5x)/(5)=-(25)/(12)*(1)/(5)
x=-(25)/(12)*(1)/(5)
x=-(5)/(12) answer

Answer by stanbon(75887)

Solve the equation
1/8 + (5/6)x = -2/9
---
Multiply thru by 72 to get:
9 + 60x = -16
60x = -25
x = -5/12
===========================
Cheers,
Stan H.

Answer by Alan3354(69443)

1/8 + 5/6x = -2/9
Multiply by 72
9 + 60x = -16
60x = -25
x = -5/12

Answer by 111456(1)